A chemist has prepared two acid solutions, one of which is 4% acid by volume, another 11% acid. HOw many cubic centimeters of each should the chemist mix together to obtain 50cm^3 of 6.24% solution?
Answer by stanbon(75887)

A chemist has prepared two acid solutions, one of which is 4% acid by volume, another 11% acid. HOw many cubic centimeters of each should the chemist mix together to obtain 50cm^3 of 6.24% solution?
---------------
Equation:
acid + acid = acid
0.04x + 0.11(50-x) = 0.0624*50
Multiply thru by 10000 to get:
400x + 1100(50-x) = 624*50
400x + 1100*50 - 1100x = 624*50
-700x = -476*50
x = 34 cm^3 (amt. of 4% solution in the mixture)
50-x = 50-34 = 16 cm^3 (amt. of 11^ solution in the mixture)
================================================================
Cheers,
Stan H.